OpenStudy (anonymous):

integrate e^(-6x)sin(e^(-3x))dx

hartnn (hartnn):

did you try substuting u = e^(-3x) ?

hartnn (hartnn):

*substituting

OpenStudy (anonymous):

yes I tried it by parts and I'm not getting the right answer.....

hartnn (hartnn):

by parts, after that substitution, right ?

OpenStudy (anonymous):

\[\int\limits e ^{-6x}\sin(e ^{-3x})dx\]

OpenStudy (anonymous):

what do you mean?

hartnn (hartnn):

note e^(-6x) = e^(-3x) * e^(-3x)

hartnn (hartnn):

i mean , you plug in u = e^-3x also find du

hartnn (hartnn):

sin e^-3x becomes sin u

OpenStudy (anonymous):

-1/3du=e^(-3x)dx

hartnn (hartnn):

yesss

hartnn (hartnn):

-1/3 integral u sin u du did u get that ?

OpenStudy (anonymous):

yes I did =)

hartnn (hartnn):

can you finish now ? let me know if you have any other doubts on this :)

OpenStudy (anonymous):

so from here I would use integration by parts?

hartnn (hartnn):

yes but if you can, i suggest you use standard result directly

OpenStudy (anonymous):

ok how would I do that ?

hartnn (hartnn):

ok, use integration by parts then

hartnn (hartnn):

or \(\int x \sin x dx = -x \cos x +\sin x + c\)

OpenStudy (anonymous):

ok thank you so much!!!

hartnn (hartnn):

welcome ^_^ and \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)